Need to solve this equation by hand. Please help me. Thanks!
2x+3 - √x+1 = 1

Answer by Edwin McCravy(20059)

2x+3 - √x+1 = 1

Isolate one of the radicals:

       √2x+3 = 1-√x+1

Square both sides

    (√2x+3)² = (1-√x+1)²
 
Squaring the left side takes away the radical
Squaring the right side means to write the 
(expression twice
    
        2x+3 = (1-√x+1)(1-√x+1)

FOIL out the right side:

        2x+3 = 1-√x+1-√x+1+(√x+1)²

Simplify

        2x+3 = 1-2√x+1+x+1
    
        2x+3 = 2+x-2√x+1

Isolate the radical term

         x+1 = -2√x+1

Square both sides:

      (x+1)² = (-2√x+1)²

  (x+1)(x+1) = (-2)²(√x+1)²


    x²+x+x+1 = 4(x+1)

     x²+2x+1 = 4x+4

Get 0 on the right:

     x²-2x-3 = 0 

Factor the left side:

  (x-3)(x+1) = 0

Us zero-factor property:

  x-3 = 0;  x+1 = 0
    x = 3;    x = -1

We must check for extraneous solutions:

Checking x = 3      

    √2x+3 - √x+1 = 1
√2(3)+3 - √(3)+1 = 1
     √6+3 - √3+1 = 1
         √9 - √4 = 1
           3 - 2 = 1
               1 = 1

therefore x = 3 is a solution:

Checking x = -1      

      √2x+3 - √x+1 = 1
 √2(-1)+3 - √(1)+1 = 1
     √-2+3 - √-1+1 = 1
           √1 - √0 = 1
             1 - 0 = 1
                 1 = 1

Therefore x = -1 is also a solution.

Edwin